OpenStudy (anonymous):

I need help with 11 6th grade questions

OpenStudy (ddcamp):

Let's try them one at a time.

sammixboo (sammixboo):

Right, always ask your questions one post at a time so we don't get confuse or spam other users

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i miss counted its less

OpenStudy (anonymous):

hi i miss counted its less than 11

OpenStudy (anonymous):

hi sorry i miss counted less than 11 questions

OpenStudy (anonymous):

:3

OpenStudy (astrophysics):

Ok so what's your first questin mr pikachu

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it was the only picture i can find good enough

OpenStudy (astrophysics):

Do you have a question?

OpenStudy (astrophysics):

That deals with math

OpenStudy (anonymous):

yes plz wait a moment

OpenStudy (anonymous):

the numbers 10,100,1000,and 10000 can be written bye the powers of 10 how can you use the prime factorization to find 270000 yes i am a 6th grade honer student

OpenStudy (anonymous):

opps i put bye instead of by sorry

OpenStudy (astrophysics):

That doesn't seem like grade 6 to me, but anyways...so \[10=10^1~~~100= 10^2~~~1000=10^3\] see the pattern here

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what u mean it dosent seem like 6th grade

OpenStudy (anonymous):

this is California curriculum they spoil us to much when i went to Chicago wow it was way harder

OpenStudy (anonymous):

I'm sorry i bragged to much that a problem i need to fix

OpenStudy (astrophysics):

Ok so the prime factors of 270,000 = 2 x 2 x 2 x 2 x 3 x 3 x 3 x 5 x 5 x 5 x 5 x 5

OpenStudy (astrophysics):

Now you just have to put it in powers

OpenStudy (astrophysics):

\[2 \times 2 \times 2 \times 2 = 2^4\] now do the others

OpenStudy (astrophysics):

Maybe this will make it more clear \[(2 \times 2 \times 2 \times 2) \times \color{red}{(3 \times 3 \times 3)} \times \color{blue}{(5 \times 5 \times 5 \times 5 \times 5)}\]

OpenStudy (astrophysics):

Oh I thought this wasn't grade 6 because when I was in grade 6 we didn't even know what powers were xD or exponents for that matter...

OpenStudy (anonymous):

thx im acually got to do the rest when you were typeing all those in and i checked them with my online tutor and she sayed they were correct thx for everyones help bye

OpenStudy (astrophysics):

But you kids are much smarter

OpenStudy (astrophysics):

Np, take care

OpenStudy (anonymous):

thx wait im not sure im trying to answer this is 2^4 x 3 ^ 2 x 5 = 720 is it because i think i did all the math

OpenStudy (astrophysics):

No the exponents are wrong, notice how I said 2 x 2 x 2 x 2 = 2^4 this means its being multiplied by itself 4 times

OpenStudy (astrophysics):

thats what the "power" of means

OpenStudy (astrophysics):

so 3 x 3 x 3 = 3^3

OpenStudy (astrophysics):

Do you see?

OpenStudy (anonymous):

oh ok

OpenStudy (astrophysics):

So 5 x 5 x 5 x 5 x 5 means?

OpenStudy (astrophysics):

Put it to the power of

OpenStudy (anonymous):

5^5

OpenStudy (astrophysics):

Yeah!

OpenStudy (astrophysics):

Simple right :)

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

ya now that you explained it to me i get it all thx for all ur help bye

OpenStudy (astrophysics):

C ya :) feel free to ask more questions anytime